OpenStudy (gabylovesyou):

Brian is creating a collage on a piece of cardboard that has an area of 110r3 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of the square root of r to the fifth power square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

OpenStudy (gabylovesyou):



OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok and the total area is \[\large 110r^3\] right?

OpenStudy (gabylovesyou):

yup

jimthompson5910 (jim_thompson5910):

well you have to solve the equation below for k \[\large k*\sqrt{r^5} = 110r^3\]

OpenStudy (gabylovesyou):

how do i solve it ?

jimthompson5910 (jim_thompson5910):

divide both sides by \(\large \sqrt{r^5} \)

jimthompson5910 (jim_thompson5910):

to isolate k

OpenStudy (gabylovesyou):

..... what would it be ? i tried a million times

jimthompson5910 (jim_thompson5910):

\[\large k*\sqrt{r^5} = 110r^3\] \[\large k = \frac{\110r^3}{\sqrt{r^5}}\] Then you rationalize the denominator

OpenStudy (gabylovesyou):

`r^3/2

jimthompson5910 (jim_thompson5910):

not quite

OpenStudy (gabylovesyou):

,-,

OpenStudy (gabylovesyou):

@jim_thompson5910 help ;P

jimthompson5910 (jim_thompson5910):

\[\large k*\sqrt{r^5} = 110r^3\] \[\large k = \frac{\110r^3}{\sqrt{r^5}}\] \[\large k = \frac{\110r^3\sqrt{r^5}}{\sqrt{r^5}\sqrt{r^5}}\] \[\large k = \frac{\110r^3\sqrt{r^5}}{(\sqrt{r^5})^2}\] \[\large k = \frac{\110r^3\sqrt{r^5}}{r^5}\] \[\large k = \frac{\110\sqrt{r^5}}{r^2}\] \[\large k = \frac{\110\sqrt{r^4*r}}{r^2}\] \[\large k = \frac{\110\sqrt{r^4}*\sqrt{r}}{r^2}\] \[\large k = \frac{\110r^2\sqrt{r}}{r^2}\] \[\large k = \110\sqrt{r}\] ------------------------------------------- So the expression \[\large \110\sqrt{r}\] models how many pieces of paper total that are needed

OpenStudy (gabylovesyou):

thats the answer? or ur asking me ?

OpenStudy (gabylovesyou):

@jim_thompson5910